OpenStudy (esraa):

An egg, initially at rest, is dropped onto a concrete surface; it breaks. Prove that the process is irreversible. In modeling this process treat the egg as the system, and assume the passage of sufficient time for the egg to return to its initial temperature

OpenStudy (esraa):

@UnkleRhaukus

OpenStudy (unklerhaukus):

the entropy

OpenStudy (esraa):

Yes, but I don't know how?

OpenStudy (unklerhaukus):

ah, energy has left the system in the form of heat

OpenStudy (esraa):

and??

OpenStudy (unklerhaukus):

and increasing the entropy

OpenStudy (unklerhaukus):

im not sure which equations you are allow to use

OpenStudy (esraa):

well the only equation we took in that chapter is \[\int\limits ds = \int\limits Cp \frac{ dT }{ T } - \int\limits \frac{ dp }{ p }\]

OpenStudy (anonymous):

The egg suffers an inelastic collision with the floor, id est energy initially in the velocity of egg's center of mass appears to vanish -- is no longer in any macroscopically-observable degrees of freedom. Since energy is conserved, it must therefore have gone into unobservable microscopic degrees of freedom, e.g. the rotation and vibrations of molecules that make up the egg. This would be observed as an increase in the temperature of the egg. However, since the egg is in thermal equilibrium with its surroundings, a temperature differential will result in a heat flow until the temperature differential is erased. We can conclude heat flows from the egg into the surroundings. Hence by the Clausius definition the entropy of the surroundings (or both the egg and the surroundings, if the surroundings are not infinite) increases. If dS > 0 then the process is irreversible.

OpenStudy (esraa):

ohhh thank u soo much 4 answering this question, but plz can u explain a bit more??!

OpenStudy (esraa):

@Carl_Pham

OpenStudy (anonymous):

Not unless you point to something that needs explaining to you.

OpenStudy (esraa):

The whole paragraph, I didn't get it!

OpenStudy (anonymous):

Sorry to hear that. Well, start reading it and thinking about it, word by word, sentence by sentence. When you come to a part on which you get stuck, ask a question. I've given you the explanation, but I'm not going to do your thinking for you, and translate it into how you would explain it to yourself, or to someone else (for example on an assignment or test). That's your job.

OpenStudy (esraa):

Well, I do appreciate that you have done the effort to answer it, but I guess you wrote it at least the first part in an improper way, so I didn't get it right!!! Ohhh and thanks a lot for being such a nice person

OpenStudy (anonymous):

Nice = doing your work for you? That's not being "nice," that's either being a patsy or helping you cheat. Either one is not nice at all, and if someone were being "nice" in that way to your student, or your own kid, you'd be angry, and quite reasonably so.

OpenStudy (esraa):

I guess you got me wrong! I wanted u to explain more and not helping me to cheat! and btw it's not my work it's an extra work, I was just curious how can I prove this by thermodynamics! And from your point of view you are absoulutely right, but I'm not cheating.......

OpenStudy (anonymous):

do you still need help with understanding this?